OpenStudy (anonymous):

A standard number cube is rolled 216 times. Predict how many times a 3 or a 5 will be the result.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

you expect a 3: 216/6 times, and a 5:216/6 times. We add these probabilities together to get the amount of times we get a 3 or 5. 216/6+216/6=432/6=72 times

OpenStudy (anonymous):

sorry not add these probabilities add these number of times together